Config13 02-Apr-2018 to 04-Apr-2018 Output Fiber: Collet Fiber on one end, FC on the other; used to replace the output fiber on Config11. Collet marked "DO NOT TWIST"; fiber labeled with orange electrical tape. This is the same fiber as Config12. Input Fiber: Collet Fiber on one end, FC on the other; used to replace the input fiber on Config12. Collet marked "DO NOT TWIST"; fiber labeled with green electrical tape. This is in line PDBS S/N 2 with the PDBS and fold mirror removed. I have redone how I connected the fibers to the PDBS. On the input side to the PDBS, I removed the Thorlabs P/N F220FC-B (aspheric lens, "B" AR coating, mounted to an FC connector) and the adaptor part that got this to the 0.5" Thorlabs lens tubes. I used lens "Vis 1 2" that I had from the S/N 1 work, which is the Clarence Z selected lenses mounted with an FC connector. I took out the FC connector from its 0.5" lens tube, where the face of the FC connector was at the inner shoulder (plus some washers) of the lens tube, and I threaded this lens tube onto the external threads of the input fiber collet. This places the tip of the fiber at the same distance as it was with the FC connector. On the output side, I undid the Config12 assembly, removed the FC connector from the "3 4 vis" lens, and threaded its lens tube (with washers) onto the external thread of the fiber collet. Following the procedure for the input fiber. The result are the input and output lenses Clarence modeled for the splitter with the fiber collet at the PDBS and FC at the input of the green fiber and output of the orange fiber. Since the fiber collet cannot be twisted, the output fiber was connected to PD2 as in Config12 -- by first disassembling the PD to FC, attaching the FC to the fiber, and threading the lens tube onto the FC to 1" part. On the input side, the FC threads into the sphere. We had put some epoxy to keep it from turning and coming out. This was removed, and all the connectors to the sphere (SEI, monitor PD, LED source) were removed so the sphere could be threaded onto the fiber. The natural position of the PDBS based on its center of gravity led to a new position for the sphere -- now the input port faces up instead of south as before. So the LED source was remounted. See the photos from April 2, 2018. The PD2 is now on gain 6. Before, it was on gain 7.